Macroeconomic

2. The IMF projects that Chinas real GDP per per- son will be 57,163 yuan in 2017 and 60,334 yuan in 2018 and that Indias real GDP per person will be 98,028 rupees in 2017 and 104,191 rupees in 2018. By maintaining their current growth rates, which country will be first to double its standard of living and when will that happen?
Answer #1


CHINA

Real GDP per person in 2017 = 57,163 Yuan

Real GDP per person in 2018 = 60,334 Yuan

Calculate the growth rate of real GDP per person -

Growth rate = [(Real GDP per person in 2018 - Real GDP per person in 2017)/Real GDP per person in 2017] * 100

Growth rate = [(60,334 - 57,163)/57,163] * 100

Growth rate = 5.55%

The growth rate of real GDP per person in China is 5.55%.

Calculate the number of years in which real GDP per person get doubled -

Number of years = 70/Growth rate of real GDP per person = 70/5.55 = 12.61 years

The China would be able to double its real GDP per person in approximately 13 years.

It means that the China would be able to double its real GDP per person by 2031.

INDIA

Real GDP per person in 2017 = 98,028 rupees

Real GDP per person in 2018 = 104,191 rupees

Calculate the growth rate of real GDP per person -

Growth rate = [(Real GDP per person in 2018 - Real GDP per person in 2017)/Real GDP per person in 2017] * 100

Growth rate = [(104,191 - 98,028)/98,028] * 100

Growth rate = 6.29%

The growth rate of real GDP per person in India is 6.29%.

Calculate the number of years in which real GDP get doubled -

Number of years = 70/Growth rate of real GDP per person = 70/6.29 = 11.13 years

The India would be able to double its real GDP per person in approximately 11 years.

It means that the India would be able to double its real GDP per person by 2029.

Thus,

By maintaining their current growth rates, India would be first to double its standard of living.

It will be able to double its standard of living by Year 2029.